OpenStudy (anonymous):

!!!!!HELP PLEASE!!! Which of the following statements is NOT true? A. Every function can be represented by a graph in the Cartesian plane. B. Another name for an x-intercept is a real zero. C. A function can have several y-intercepts. D. A function can have infinitely many x-intercepts.

OpenStudy (anonymous):

i would say both A and C

OpenStudy (anonymous):

I can only choose one though.

OpenStudy (anonymous):

but you are supposed to say C since if it has more then one y intercept it clearly does not pass the "vertical line test" since the y axis is a vertical line who wrote this?

OpenStudy (anonymous):

ask who ever wrote it to plot the function defined by \[ f(x) = \left\{ \begin{array}{lr} 1 & : x \in \mathbb{Q}\\ 0 & : x \notin \mathbb{Q} \end{array} \right.\]

OpenStudy (anonymous):

It is an online assignment kinda thing so I can't ask but thank you for your help C was correct.

OpenStudy (anonymous):

then tell them they need to learn some math ok, not if it is your math teacher...

OpenStudy (anonymous):

let me guess FVLA or sommat?

OpenStudy (anonymous):

My Math Lab is what it is but yeah something like that.

OpenStudy (anonymous):

pearson should know better